OpenStudy (anonymous):

please help me solve 1 problem with 3 variables and 3 equations. 11.) 2x+3y+2z=1 x+4y-z=7 3x+y+3z=-2 elimination please. what do I do first??

OpenStudy (agent0smith):

Substitution would be easier... do you have to use elimination? to eliminate z in the first equation: 2x+3y+2z=1 +2(x+4y-z=7) ---------------

OpenStudy (anonymous):

no we don't have to use elimination I find it easier to understand though. it says substitution orr elimination.. please explain what u put above?

OpenStudy (anonymous):

I got 4x+11y=8 using it your way sir? @argent0smith ??

OpenStudy (agent0smith):

I'd recommend substition then (it's easier with 3 equations. 2 equations elimination is easier). You did that right (except 8 should be 15), but elimination is...ugh with three. Solve this equation for x =.... x+4y-z=7

OpenStudy (anonymous):

im lost so badly. ???? please help me with www.vyew.com and post link on here

OpenStudy (agent0smith):

Just rearrange this equation until it looks like x= x+4y-z=7 start by subtracting 4y on both sides

OpenStudy (anonymous):

4x+11y=15 -> 11y+4x=15 x+4y-z=7-> 4y-z=7 is that right?

OpenStudy (anonymous):

im LOSTtttt??????

OpenStudy (agent0smith):

Just use x+4y-z=7, we aren't using the other equation. subtracting 4y on both sides gives x-z=7 - 4y now add z to both sides

OpenStudy (dumbcow):

it all depends on your preference, i find elimination to be easier even for 3 equations

OpenStudy (anonymous):

now what /: elimination I understand better I don't understand his method at allll

OpenStudy (dumbcow):

either way you have to eliminate a variable to turn it into a system of 2 equations, then solve that system

OpenStudy (agent0smith):

Okay then go with what we have above: 4x+11y=15. Leave that alone for now, don't use it. now do this one: 3(x+4y-z=7) -3x+y+3z=-2 ------------------

OpenStudy (agent0smith):

(that's the middle and last equations)

OpenStudy (agent0smith):

oops, 3(x+4y-z=7) -(3x+y+3z=-2) ------------------

OpenStudy (agent0smith):

That'll actually allow us to solve for y...

OpenStudy (anonymous):

im losttt @agent0smith please help me the elimination way

OpenStudy (agent0smith):

That is the elimination way.

OpenStudy (anonymous):

I got this 3x+4y-z=7 -3x+y+3z=2 = 5y-2z=9 Right?????

OpenStudy (agent0smith):

3(x+4y-z=7) -(3x+y+3z=-2) ------------------ the things outside the brackets apply to EVERY term inside the brackets. Not just the first term.

OpenStudy (anonymous):

12y+6z=7 ??

OpenStudy (agent0smith):

actually we need to add them, to eliminate z, not x 3(x+4y-z=7) +(3x+y+3z=-2) ----------------- first distribute the 3 outside the brackets 3x+12y-3z=21 (multiply all by 3) +(3x+y+3z=-2) -----------------

OpenStudy (anonymous):

I did multiply everything by 3 and I added what I got and it come to 12y+6z=7

OpenStudy (agent0smith):

Use the correction above. 3x+12y-3z=21 +(3x+y+3z=-2)

OpenStudy (anonymous):

I did already... U keep repeating urself but im still Lost..sir?

OpenStudy (agent0smith):

12y+6z=7 is not correct, for the above

OpenStudy (anonymous):

I added and multiplied by 3 what u put and got 12x+ 37y-6z=19 for a total after I multiplied and added

OpenStudy (agent0smith):

That's not right. Add together all the terms that are lined up vertically. 3x +12y -3z = 21 + 3x +y +3z = -2

OpenStudy (anonymous):

6x+13y=19 ? and howd u get those 2 equations?? still lost )":

OpenStudy (agent0smith):

We got them from above, scroll way back up. And 6x+13y=19 is correct. Now you need to go back to this one from earlier... 4x+11y=15 and solve them together: 6x+13y=19 4x+11y=15

OpenStudy (anonymous):

my laptop is jetlaging I got 10x+24y=34

OpenStudy (anonymous):

now what??????

OpenStudy (agent0smith):

We can't just add them together. You need to solve them using elimination. We can eliminate x like so 2(6x+13y=19) -3(4x+11y=15) -------------- don't forget to apply the numbers outside to everything in the brackets

OpenStudy (anonymous):

y=1 ????? /:

OpenStudy (agent0smith):

Yes. Now use any equation to solve for x, we know what y is... like this one 4x+11y=15

OpenStudy (anonymous):

thanks I got it from here. im still lost until I get to here tho would u help me set up 12. just tell me what to do??? like put it step by step please??

OpenStudy (anonymous):

12.) x-2y+3z=-7 4x+5y+z=4 -x+y-2z=5

OpenStudy (agent0smith):

Start with the first and last, since you can eliminate x easily add them together: x-2y+3z=-7 -x+y-2z=5 then add the middle and last, like so, to eliminate x 4x+5y+z=4 +4(-x+y-2z=5) then, you'll have two equations from those with only y and z, which you can solve together.

OpenStudy (anonymous):

I still have x on one of them? I got -y+ -z=12 and I got -3x+6y-z=9

OpenStudy (agent0smith):

Check your work, 4x minus 4x is zero. -y+ -z=12 is not correct... z should be positive, -7 plus 5 is not 12.

OpenStudy (anonymous):

ok I redid it and got.. 6y-z=9 + -y+-z=12

OpenStudy (agent0smith):

The first, adding them: x-2y+3z=-7 + -x+y-2z=5 ----------- -y + z = -2 <=== that is what you should get

OpenStudy (anonymous):

now what?

OpenStudy (agent0smith):

Have you done this one correctly yet? the -3x+6y-z=9 you posted isn't right (adding them) 4x+5y+z=4 +4(-x+y-2z=5)

OpenStudy (anonymous):

yeah I got -y+z=-2 but idk how u got that??

OpenStudy (agent0smith):

how i got what?

OpenStudy (anonymous):

how u got -y+z=-2

OpenStudy (anonymous):

I got 6y-z=9 im lost? is that right.??

OpenStudy (agent0smith):

I got it by adding the two together. That isn't right. 4x+5y+z=4 +4(-x+y-2z=5) 4x+5y+z=4 + -4x+4y-8z=40 (after multiplying all of it by 4)

OpenStudy (anonymous):

9y+-7z=44

OpenStudy (agent0smith):

Yes now solve these two (you can't just add them as is, use elimination) -y + z = -2 9y-7z=44